hen x grams of carbon are heated with y grams of oxygeb in a closed vessel, no residue is left behind.. Which of the following statements is correct? A) y/x must lie between 1. 33 & 2. 67B) y/x must be greater than or equal to 2. 67C) y/x must be less than or equal to 1. 33D) y/x must be greater than or equal to 1. 33

Grade:11

1 Answers

varun
6 years ago
HELLO KIRTI LODHA
understand this,
If one mole of each reactant is used in the reaction then,
C + O2 gives CO2
12gm  32 gm     44 gm
(X)       (Y)
therefore,, 
according to the question,if no residue is left behind then CO2 will escape as a gas
SO,,
X/Y=12/32
      =0.375
AND
Y/X=2.6667 ======= which is greater number
which is approximately equal to 2.67
SO OPTION B IS CORRECT OPTION
THANK YOU